OpenStudy (anonymous):

Can you approve of my calculation please? my question and answer are below and its about vectors

ganeshie8 (ganeshie8):

looks legit

ganeshie8 (ganeshie8):

another method :- say, \(\theta\) is the angle between \(\vec{b}\) and \(\vec{a}\), then \(-\theta\) will be the angle between \(\vec{a}\) and \(\vec{b}\), \(\large \vec{b} \times \vec{a} = |a||b|\sin (\theta) \hat{N}\) \(\large \vec{a} \times \vec{b} = |a||b|\sin (- \theta) \hat{N} = -\vec{b} \times \vec{a} \)

OpenStudy (anonymous):

so it could be written as simple as that?

ganeshie8 (ganeshie8):

yup ! i like the second method better

OpenStudy (anonymous):

i don't think we studied this formula yet haha .. but how would i improve my method .. that could be better haha

ganeshie8 (ganeshie8):

it is a crime to study scalar triple products, w/o knowing this definition for vector product. i saw u doing scalar triple product problems in ur previous questions :o

OpenStudy (anonymous):

oh then maybe we did but i didn't pick it up ill will go over our thing again haha

ganeshie8 (ganeshie8):

okay ! but ur method is completely legit.

OpenStudy (anonymous):

oh ok so there were no mistakes? if that then i will use that one

ganeshie8 (ganeshie8):

i had a cursory glance, it is correct. g

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

btw is it better to submit questions with answers in this forum or just the question?

ganeshie8 (ganeshie8):

you can simply post the question, and start working it here if u want. thats one way. if you had worked it already, just take a screenshot and post it and discuss. if u are stuck on something... then u can simply start the question wid, "I am stuck on this... " :)

ganeshie8 (ganeshie8):

you dont have to post ur work everytime.... but its good to show ur work when u can - that way others know wat exactly u are working on, right...

OpenStudy (anonymous):

ok thanks